Average inventory = 10,000 units Peak inventory (April) = 24,000 units Low inventory (October) = 5,000...

  • Average inventory = 10,000 units
  • Peak inventory (April) = 24,000 units
  • Low inventory (October) = 5,000 units
  • Storage = 2%
  • Taxes = 1%
  • Inflation = 4%
  • Risk = 1.5%
  • Pilferage = .5%
  • Service = 2.5%
  • Cost of capital = 20%
  • Prime lending rate = 3.5%
  • Insurance = 1.5%
  • Cost of one unit = $5.00
  • Lead-time = 5 days
  • Order cost = $250

What is the annual inventory carrying cost for the War Eagle Brand golf ball, to the nearest whole dollar? Do not enter a dollar sign in your answer.
Answer #1

Annual inventory Carrying Cost = Average Inventory*Carrying cost

Average inventory = 10,000

Carrying cost = Storage*Cost of one unit = 5*0.02 = $0.1

Annual inventory Carrying Cost = 10,000*0.1 = 1,000